OpenStudy (anonymous):

The graph of y=(5x^4)-x^5 has inflection points where?

OpenStudy (anonymous):

The answer choices are: (A) (0,0) only (B) (3,162) only (C) (4,256) only (D) (0,0) and (3,162) (E) (0,0) and (4,256)

OpenStudy (anonymous):

I thought it was only (3,162) because the second derivative only changes signs at x=3, but the graph shows that (0,0) is also a point of inflection.

OpenStudy (zale101):

\(\large f(x)=5x^4-x^5 \) \(\large {f'(x)=20x^3-5x^4}\) \(\large f''(x)=60x^2-20x^3\) set the second derivative equal to zero and test points between those given x's. If there's a change in concavity then that means there's in inflection point happening in the graph. \(\large {-20x^2(x-3)=0}\) \(\large x=3\) and \(\large x=0\)